OpenStudy (anonymous):

a photograph measures 6in by 9in. earl put in a frame that added 2 in to every side. find the outside perimeter and total area of the photograph and frame

OpenStudy (anonymous):

photograph area = 9*6 obviously. perimeter = 2*6+2*9 frame area = (6+4)*(9+4). as you add 4 to width and height. perimeter = 2*(6+2) + 2*(9+2) get it?

OpenStudy (anonymous):

sorta

OpenStudy (campbell_st):

 here is the picture